OpenStudy (anonymous):

find the areas of increasing and decreasing and extremum points of the function: y=8X2-Inx

OpenStudy (anonymous):

y=x-sinx

OpenStudy (anonymous):

can anybody help me solve the above two questions

OpenStudy (anonymous):

i can help if you rewrite the first one is it \[f(x)=8x^2-\ln(x)\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok so \[f'(x)=16x-\frac{1}{x}\] and your job is to see where this is positive and negative, that is solve \[16x-\frac{1}{x}>0\]

OpenStudy (anonymous):

f(x)=2x3-15x2+36x-2

OpenStudy (anonymous):

\[\frac{16x^2-1}{x}>0\] \[\frac{16(x+\frac{1}{4})(x-\frac{1}{4})}{x}>0\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

so you have 4 intervals to consider \[(\infty, -\frac{1}{4}),(-\frac{1}{4},0),(0,\frac{1}{4}), (\frac{1}{4},\infty)\] and it will be positive on the second and 4th, negative of the first and third, so that gives intervals of increase and decrease respectively

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

same idea for \[f(x)=x-\sin(x)\] \[f'(x)=1-\cos(x)\]

OpenStudy (anonymous):

what of f(x)=2x3-15x2+36x-2

OpenStudy (anonymous):

same idea

OpenStudy (anonymous):

we can do that next, but notice that \[1-\cos(x)>0\iff 1>\cos(x)\] so that one is always increasing

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

\[f(x)=2x^3-15x^2+36x-2\] \[f'(x)=6x^2-30x+36\] so this one is a quadratic and we can find the zeros probably by factoring

OpenStudy (anonymous):

\[x^2-5x+6>0\] \[(x-3)(x-2)>0\] and since this thing is a parabola that opens up it will be negative between the zeros and positive outside them, so your cubic polynomial will be increasing on \[(-\infty, 2)\cup (3,\infty)\] and decreasing between

OpenStudy (anonymous):

find the velocity and acceleration of the process: x=10e-2t.cos (5t + π/2)